Ca = 1500 – 20 r c = 0.6 G = T = 2000 Ip = 2500 – 40 r NX = - 300 Derive Ap as a function of interest rate Calculate general multiplier Derive planned expenditure Ep as a function of interest rate If r = 5, calculate the equilibrium income If Y = 11000 and r = 5, what is planned investment? What is actual investment? What is unplanned investment or inventory change?
Answer #1

Given that in o Ap = ca tlp tot NX Ap = 1500 +2500-gor + 2000-300 Ap = 5700 407 (6) multiplico 1/(1-C) =1/11 - 0.06) = 104 -3 (d) where 8=5, Ep = 4500 tooby -4085 EP = 4500 +0.64 - 200 EP: 4300 +0.69 setting Y=EP Y = 4300 +0.6 Y 0-4 Y = 4300 = 10,75